OpenStudy (anonymous):

Jose received $30 for his birthday that he put in his piggy bank. Each week, he puts 4 more dollars in his bank so that the amount of money in dollars D in his bank is given by D = 4w + 30 where w is the number of weeks. His sister Marisol makes a table of reasonable values for w to determine how much money Andre will eventually have. If one of the values for w Marisol selects is 36, what would be her corresponding entry for D? 466 174 70 39

OpenStudy (emmamink):

It is 174 because if you plug in 36 for w in the equation, that is what you would get.

OpenStudy (michele_laino):

please try to substitute w=36 in your formula D=4*w+30.

OpenStudy (anonymous):

i got 174 also

OpenStudy (emmamink):

D = 4(36) + 30. D = 144 + 30. D = 174

OpenStudy (emmamink):

Please, medal if you appreciate my help and tag me anytime.

OpenStudy (anonymous):

Thank you both

OpenStudy (emmamink):

Sorry. I have to go.

OpenStudy (anonymous):

k bye

OpenStudy (anonymous):

thanks for your help

OpenStudy (anonymous):

@Michele_Laino can you help

OpenStudy (michele_laino):

ok! I try...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i figured it out

OpenStudy (michele_laino):

please, I conjeture the subsequent linear relationship, between D and w: \[D=a*w+b\] where a ad b are o coefficent to be determined as follows:

OpenStudy (michele_laino):

if w=4, we have D=58, so I can write: (1) \[58=4a+b\] wheres if w=15, then D=80, so we can write: (2) \[80=15a+b\] Now, solving the equation (1), wean write: (3)\[b=58-4a\] inserting quation (3) in equation (2), we have: \[80=15a+58-4a\] from which I get: \[a=2\] so your anser is the second amon those listed @welligya...